OpenStudy (anonymous):

Evaluate the expression for a = –1 and b = 5. (ab)to the second power

OpenStudy (anonymous):

Does it equal -25?

OpenStudy (anonymous):

Substitute the given into the parenthesis. (-1x5)^2=(-5)^2 = 25...since the negative is in the parenthesis.

OpenStudy (anonymous):

Oh, So it is positive 25 then? Why is that?

OpenStudy (anonymous):

Sorry I mistyped

OpenStudy (anonymous):

It's -25

OpenStudy (anonymous):

\[(ab)^{2}=a ^{2}b ^{2}\] It will be positive.

OpenStudy (anonymous):

@Kuromeru, @sunshnee8 , Why is it positive?

OpenStudy (anonymous):

Isn't the rule...multiply what's inside first before you distribute the exponent?

OpenStudy (anonymous):

@sunshnee8- My calculator says -25... I did -1x5, and then to the second power

OpenStudy (anonymous):

Yes, if you did the answer to the second power the calculator does not automatically keep the ( ) so therefore it will yield a negative answer. If you put it in manually it should give you a positive answer.

OpenStudy (anonymous):

Any negative number to the power of a positive number is positive.

OpenStudy (anonymous):

@sunshnee8 It is, but if you look at (ab)^2, it = ab(ab) or a^2b^2